OpenStudy (anonymous):

You draw 2 cards from a well-shuffled deck of 52 cards. After you draw the first card, you do not replace it in the deck. What is the probability of P(get ace of spades on first draw and get queen of hearts on second draw)?

OpenStudy (anonymous):

1/52

OpenStudy (anonymous):

1/52*1/51

OpenStudy (anonymous):

1/51

OpenStudy (anonymous):

andras is correct. \[P(A\cap B)=P(A)P(B|A)\]

OpenStudy (anonymous):

I will have applied probability next year, cant wait!

OpenStudy (anonymous):

my answer would be two fractions then if I have to keep the answer in a fraction?

OpenStudy (anonymous):

A = the event the first card is the ace of spades, B is the event the second card is the queen of hearts \[P(A)=\frac{1}{52}\] \[P(B|A)=\frac{1}{51}\]

OpenStudy (anonymous):

product is \[\frac{1}{2652}\]

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

Andras have fun! it is a great topic.